# Flatten Long-Range Loss Landscapes for Cross-Domain Few-Shot Learning

### v1

@@ 1 Introduction
To handle this problem, we notice that we can easily find different minima in the representation space: different normalization methods can produce different representations, all of which remain suitable for classification. Since the distance between representations is much larger than the sunken region in the loss landscape around each minimum, we can flatten a long-range loss landscape between these minima (i.e., differently normalized representations).
Based on the above intuition, we randomly interpolate these differently normalized representations at each layer of deep networks, and utilize interpolated representations for classification. This would push the interpolated representations to be effective, which lowers the high-loss region between minima and flattens the loss landscape. Specifically, we provide two instantiations for both Convolutional Neural Networks and Vision Transformers, and implement the above operations as a normalization layer (FLoR layer) to replace the original normalization layer.
To sum up, our contribution can be listed as
@@ 2.1 Preliminaries
where $F(x_{i},y_{i})=h(f(x_{i}))$ outputs the classification probability of $x_{i}$, which is composed of a feature extractor $f(\cdot)$ and a classifier $h(\cdot)$. Then, $f(\cdot)$ is transferred to novel-class dataset $D^{novel}=\{x^{u}_{i},y^{u}_{i}\}_{i}^{N^{u}}$ where $y^{u}_{i}\in C^{novel}$ (novel-class stage). Typically, only 1 or 5 training samples are available for each novel class, which makes the training on novel classes difficult. For a fair comparison, during the novel-class stage, current works [11] always sample $k$-way $n$-shot episodes for novel-class training and evaluation. Each episode contains a support set $S=\{x^{s}_{ij},y^{s}_{ij}\}_{i,j}^{k,n}$ with $k$ novel classes and $n$ training samples in each class for training, and a query set $Q=\{x^{q}_{i}\}_{i}^{N^{q}}$ for evaluation. Denote the model trained on novel classes as $F^{u}(\cdot)=h^{u}(f^{u}(\cdot))$, the prediction for $x^{q}_{i}$ is
@@ 2.2 Representation-Space Loss Landscape
Figure 2: To validate the analysis based on representation-space loss landscapes (RSLL), we apply low-frequency noises to the representation space (i.e., pixels and features) to be the domain shifts on training data for a sanity check. The perturbation variance measures the distance between the perturbed representation and the original representation (a minimum in RSLL). We use the performance drop against perturbation variance to measure the sharpness of the landscapes around the minimum, where a larger drop indicates a sharp minimum. We can see the model based on Instance Normalization (IN) is located in a flatter minimum than the model based on Batch Normalization (BN), which brings the high performance of the IN-based model in cross-domain tasks. This result is consistent with current works [10] and validates the rationale of the RSLL analysis. (a) Samples of the pixel perturbation. (b) Perturbation on pixels. (c) Perturbation on features.
@@ 2.3 Analyzing Model Generalization
We first apply pixel perturbation (i.e., representing each data point by RGB pixels) to the training data. Current works [9] show that the domain information is always correlated with the low-frequency information in the Fourier-transformed images. Therefore, we randomly sample a low-frequency perturbation from the Gaussian Distribution, where the Variance controls the distance between the original image and the perturbed image. The perturbation is at the shape of $h^{p}\times w^{p}$ where $h^{p}$ and $w^{p}$ are much smaller than the image size to ensure the frequency is low. Then, the perturbation will be resized to the size of each image and finally applied to them. Examples of such perturbation are shown in Fig. 2a, and the magnitude of the performance drop is shown in Fig. 2b. As can be seen, the performance of all models drops under the perturbations, but the IN model drops less than the BN model, indicating its robustness to the domain shift. This result is consistent with current works [10], and verifies the IN model makes training images locate in a flatter minimum in the loss landscape.
@@
Then, we apply the perturbation to the feature maps (i.e., represent each sample in the feature space). The perturbation is also the low-frequency Gaussian noise, and the magnitude of performance drop is in Fig. 2c, where we can see similar results as in Fig. 2b. This result verifies the RSLL analysis also fits the feature map representation for each input sample, validating the rationale.
To verify this analysis applies to real-world domain shifts, we then construct 5 domain-shifted miniImageNet by selecting the same classes from ImageNet-c [14]. The domain shifts contain blur, digital, extra, noise, and weather perturbations. Examples and results can be found in Tab. 1. This dataset provides 5 levels of perturbations for each kind of perturbation, therefore we utilize this level to represent the distance between the perturbed representation and the original one (a minimum in the RSLL). Similar to Fig. 2, the sharpness is measured by the performance drop. We can see the results are consistent with the toy experiments in Fig. 2, which further validates the rationale of the RSLL analysis.
@@
Table 2: Our perturbation step size (flattened range) is much larger than SAM in each representation, no matter when it is at the beginning (Init.) or the end of the training.
@@
In summary, by comparing IN and BN models on domain-shifted datasets, experiments in Fig. 2 and Tab. 1 validate (1) the RSLL analysis is rationale; (2) the domain perturbation on pixels and features, which we jointly named as representation, are consistent. Moreover, it also interprets why IN models are better, as IN models’ representations are in flatter minima compared with BN models.
@@ 2.4 Short-Range Flatness Limits Generalization
Moreover, we can observe the step sizes are small, but the IN-based model could tolerate a larger step size with a higher performance. Combining this result with the results in Fig. 2 and Tab. 1, we hold that this is because the IN-based model extracts features in a flatter RSLL minimum. As the perturbation direction is obtained by $\nabla f(x_{i})$ which only estimates the local shape of RSLL, when the minimum is sharp, it fails to estimate the complex loss landscape shape outside the local minimum of $f(x_{i})$. Therefore, the perturbation direction $\nabla f(x_{i})$ will be less effective when the step size is large. In contrast, since IN-based representations are in flatter minima than BN-based representations, it ensures $\nabla f(x_{i})$ is still effective given a larger perturbation step size. In this scenario, a larger step size would help the model to flat a longer range of RSLL, benefiting model generalization.
@@ 3 Flatten Long-Range Loss Landscapes
To benefit the cross-domain few-shot learning (CDFSL) task with improved transferability and finetuning, we aim to flatten the representation-space loss landscapes for the source-domain training. To enlarge the range of flattened area, basically rely on the local information (such as gradient [8] or curvature [34]) of the optimum may not be enough, because this information is only effective inside the sunken area in the landscape.
To handle this problem, we notice a distinguished difference between the PSLL and RSLL: there are many easy-to-find minima in RSLL. For example, previous works proposed many different normalization methods, such as Batch Normalization or Instance Normalization. This method can provide effective representations that are (almost) comparable to each other, although they may show differences in the fitness for different tasks. In other words, given the same training sample, different normalization methods could provide different minima in the same representation-space loss landscape. This inspires us to flatten the loss in the area between these minima (Fig. 1c).
@@
where $f_{norm1}(x)$ and $f_{norm2}(x)$ refer to two normalized representations, and $\delta\in[0,1]$ is an interpolation parameter randomly sampled from the Beta distribution. In implementation, we carry out the above interpolation in every layer’s representation. Since the classification is based on the interpolated representation, this representation will be pushed to be effective, so that it will be mapped to a low loss in the RSLL. Therefore, the high-loss region between $f_{norm1}(x)$ and $f_{norm2}(x)$ will be flattened.
Since this method is agnostic to the shape of the complex loss landscape between these minima, we do not need to consider any local information around minima. We measure the distance between the BN and IN representations in Tab. 2, and we can see the distance of each representation is much larger than the step size in Fig. 3, indicating a much larger flattened region.
@@ 4 Experiments
Experiments are conducted on 8 cross-domain few-shot learning (CDFSL) datasets, and our code will be released.
@@ 4.3 Comparison with state-of-the-art works
Comparisons with works utilizing ResNet10 are listed in Tab. 4 and 5 for 1-shot and 5-shot scenarios respectively. For fairness, we group work by whether finetuning (FT) or the transductive (TR) setting is used, and compare with each group separately. For all groups, we follow MN to use the cosine similarity, and subtract the mean of the support (query) set for inductive (transductive) calibration. For the transductive setting, we pseudo-label the query set by the prototype-based classifier or the finetune-based classifier, and use the pseudo-labeled query set to further update the classifier until convergence. As can be seen from Tab. 4 and 5, we can achieve the top average performance in all settings, and achieve the highest performance in almost all datasets. Notably, on ISIC2018, we can significantly outperform state-of-the-art works by as much as 9% in the 5-shot setting. Comparisons with works using ViT-S are in Tab. 6, and we also achieve state-of-the-art performance.
@@ 4.4.2 Verification of each design
Table 7: Ablation study of Mixup, IN, perturbations and $\delta$ by 5-way 5-shot accuracy with plain prototype classifier.

### v2

@@ 1 Introduction
To handle this problem, we notice that we can easily find different minima in the representation space: given an un-normalized input representation, many normalization methods have been proved to be effective in producing discriminative normalized outputs, which can be viewed as different minima in the representation space. As the distance between representations is much larger than the sunken region in the loss landscape around each minimum, we can flatten a long-range loss landscape between these minima.
Based on this intuition, we randomly interpolate differently normalized representations at each layer of deep networks, and use interpolated representations for classification. This would push interpolated representations to be effective, lowering the high-loss region between minima and flattening the loss landscape. Specifically, we provide two instantiations for both Convolutional Neural Networks and Vision Transformers, and implement the above operations as a normalization layer (FLoR layer) to replace the original normalization layer. Our contribution can be listed as
@@ 2.1 Preliminaries
where $F(x_{i})=h(f(x_{i}))$ outputs the classification probability of $x_{i}$, which is composed of a feature extractor $f(\cdot)$ and a classifier $h(\cdot)$. Then, $f(\cdot)$ is transferred to novel-class dataset $D^{novel}=\{x^{u}_{i},y^{u}_{i}\}_{i}^{N^{u}}$ where $y^{u}_{i}\in C^{novel}$ (novel-class stage). Typically, only 1 or 5 training samples are available for each novel class, which makes the training on novel classes difficult. For a fair comparison, during the novel-class stage, current works [12] always sample $k$-way $n$-shot episodes for novel-class training and evaluation. Each episode contains a support set $S=\{x^{s}_{ij},y^{s}_{ij}\}_{i,j}^{k,n}$ with $k$ novel classes and $n$ training samples in each class for training, and a query set $Q=\{x^{q}_{i}\}_{i}^{N^{q}}$ for evaluation. Denote the model trained on novel classes as $F^{u}(\cdot)=h^{u}(f^{u}(\cdot))$, the prediction for $x^{q}_{i}$ is
@@ 2.2 Representation-Space Loss Landscape
Figure 2: To validate the analysis based on representation-space loss landscapes (RSLL), we apply low-frequency noises to the representation space (i.e., pixels and features) to be the domain shifts on training data for a sanity check. The perturbation variance measures the distance between the perturbed representation and the original representation (a minimum in RSLL). We use the performance drop against perturbation variance to measure the sharpness of the landscapes around the minimum, where a larger drop indicates a sharper minimum. We can see the model based on Instance Normalization (IN) is located in a flatter minimum than the model based on Batch Normalization (BN), which brings the high performance of the IN-based model in cross-domain tasks. This result is consistent with current works [11] and validates the rationale of the RSLL analysis. (a) Samples of the pixel perturbation. (b) Perturbation on pixels. (c) Perturbation on features.
@@ 2.3 Verification and Interpretation
We first apply pixel perturbation (i.e., representing each data point by RGB pixels) to the training data. Current works [10] show that the domain information is always correlated with the low-frequency information in the Fourier-transformed images. Therefore, we randomly sample a low-frequency perturbation from the Gaussian Distribution, where the Variance controls the distance between the original image and the perturbed image. The perturbation is at the shape of $h^{p}\times w^{p}$ where $h^{p}$ and $w^{p}$ are much smaller than the image size to ensure the frequency is low. Then, the perturbation will be resized to the size of each image and finally applied to them. Examples of such perturbation are shown in Fig. 2a, and the magnitude of the performance drop is shown in Fig. 2b. As can be seen, the performance of all models drops under the perturbations, but the IN model drops less than the BN model, indicating its robustness to the domain shift. This result is consistent with current works [11], and verifies the IN model makes training images located in a flatter minimum in RSLL
@@
Then, we apply perturbations to feature maps (i.e., represent each sample in the feature space). The perturbation is also the low-frequency Gaussian noise, and the performance drop is in Fig. 2c, where we can see similar results as in Fig. 2b. This result verifies the RSLL analysis also fits the feature map representation for each input sample.
To verify this analysis applies to real-world domain shifts, we then construct 5 domain-shifted miniImageNet by selecting the same classes from ImageNet-c [15]. The domain shifts contain blur, digital, extra, noise, and weather perturbations. Examples and results can be found in Tab. 1. This dataset provides 5 levels of perturbations for each kind of perturbation, so we utilize this level to represent the distance between the perturbed representation and the original one (a minimum in the RSLL). Similar to Fig. 2, the sharpness is measured by the performance drop. We can see the results are consistent with the toy experiments in Fig. 2, which further validates the rationale of the RSLL analysis.
@@
Table 2: Our perturbation step size (flattened range) is much larger than SAM in the representation output by each layer, no matter when it is at the beginning (Init.) or the end of the training.
@@
In summary, by comparing IN and BN models on domain-shifted datasets, experiments in Fig. 2 and Tab. 1 validate (1) the RSLL analysis is rationale; (2) the domain perturbation on pixels and features, which we uniformly name as representation, are consistent. Moreover, it also interprets why IN models are better, as IN models’ representations are in flatter minima compared with BN models.
@@ 2.4 Short-Range Flatness Limits Generalization
Moreover, we can observe the step sizes are small, but the IN-based model could tolerate a larger step size with a higher performance. Combining this result with the results in Fig. 2 and Tab. 1, we hold that this is because the IN-based model extracts features in a flatter RSLL minimum. As the perturbation direction is obtained by $\nabla f(x_{i})$ which only estimates the local shape of RSLL, when the minimum is sharp, it fails to estimate the complex loss landscape shape outside the local minimum of $f(x_{i})$. Therefore, the perturbation direction $\nabla f(x_{i})$ will be less effective when the step size is large. In contrast, since IN-based representations are in flatter minima than BN-based ones, it ensures $\nabla f(x_{i})$ is still effective given a larger perturbation step size. In this scenario, a larger step size would help the model to flat a longer range of RSLL and improve generalization.
@@ 3 Flatten Long-Range Loss Landscapes
To enlarge the range of flattened area, basically relying on the local information (such as gradient [9] or curvature [37]) of the optimum may not be enough, because this information is only effective inside the sunken area in the landscape.
To handle this problem, we notice a distinguished difference between the PSLL and RSLL: there are many easy-to-find minima in RSLL. For example, previous works proposed many different normalization methods, such as Batch Normalization or Instance Normalization. These methods can provide effective representations suitable for classification, although they may show differences in the fitness for different tasks. In other words, given the same training sample, different normalization methods could provide different minima in the same representation-space loss landscape. This inspires us to flatten the loss in the area between these minima (Fig. 1b).
@@
where $f_{norm1}(x)$ and $f_{norm2}(x)$ refer to two normalized representations, and $\delta\in[0,1]$ is a ratio randomly sampled from the Beta distribution. In implementation, we conduct the interpolation in every layer’s representation. Since the classification is based on the interpolated representation, this representation will be pushed to be effective, so it will be mapped to a low loss in the RSLL. Thus, the high-loss region between $f_{norm1}(x)$ and $f_{norm2}(x)$ will be flattened.
As this method is agnostic to the shape of the complex loss landscape between minima, we do not need to consider any local information around minima. We measure the distance between the BN and IN representations in Tab. 2, where the distance of each representation is much larger than step sizes in Fig. 3, indicating a larger flattened region.
@@ 4.3 Comparison with state-of-the-art works
Comparisons utilizing ResNet10 are listed in Tab. 4 and 5. For fairness, we group work by whether finetuning (FT) or the transductive (TR) setting is used. For all groups, we follow MN to use the cosine similarity, and subtract the mean of the support (query) set for inductive (transductive) calibration. For the transductive setting, we pseudo-label the query set by the prototype-based classifier or the finetune-based classifier, and use the pseudo-labeled query set to further update the classifier until convergence. As shown in Tab. 4 and 5, we achieve the top average performance in all settings, and achieve the highest performance in almost all datasets. Notably, on ISIC2018, we can significantly outperform state-of-the-art works by as much as 9% in the 5-shot setting. Comparisons with works using ViT-S are in Tab. 6, and we also achieve state-of-the-art performance.
@@ 4.4.2 Verification of each design
Table 7: Ablation study of RSLL, Mixup, IN, perturbations and $\delta$ by 5-way 5-shot accuracy with plain prototype classifier.
@@ 6 Conclusion
Acknowledgments
This work is supported by National Natural Science Foundation of China under grants 62206102, 62376103, 62302184, and Science and Technology Support Program of Hubei Province under grant 2022BAA046.